Ticket Number: HEA-NOC/20080728-1 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: BT Ireland
Ticket Opened: 20080728 10:53 UTC+1 Problem Start: 20080728 10:31 UTC+1
Ticket Closed: 20080806 12:04 UTC+1 Problem End: 20080730 09:45 UTC+1
NCAD lost connectivity for approx 8 minutes.
Connecivity to and from NCAD
20080728 10:51 KD Contacted BT, and because the line is now back up they
are contacting networking centre to schedule a priority 3 error check on
the line. REF 1-595254031
20080728 11:18 KD Second flap logged 11:07-11:12
20080728 11:48 KD BT have responded saying there was an issue with a Laser
in Fenian st, but it is not stable. They are leaving the ticket open for
the next few hours to keep an eye on things.
20080729 14:36 KD Connection flap again at today
14:30, contacting BT.
20080730 09:51 KD 3rd day having a flap 9:22am-9:45am. Contacting BT BT.
Scheduled for alarm check. Raising the priority of the check.
20080730 11:40 GAR Call from Peter in BT. He has asked the SDH engineers
to take a look and they say they are seeing errors on the Fenian St. end.
The suspicion is that there is some bad/loose cabling at this end.
Seeing AIS and remote alarms, suggesting that it's loose cabling on our
Will request that the cabling be checked on both sites and respond to BT.
BT will hold the ticket open while this cabling is checked.
20080731 12:06 GAR Contacted NCAD and requested that they check for any
loose fibre patch leads.
20080731 12:46 GAR All fibre looks secure, both ends. Passed back to BT
and will continue to monitor for 24 hours to see if a failure recurs.
20080806 12:04 GAR Failure has not been seen again. Closing.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080728-1
HEAnet Limited. Registered in Ireland, No. 275301.